log(base 3)(x-5)+ log(base 3)(x+3)=2
Answer by scott8148(6628)

(x - 5)(x + 3) = 9

x^2 - 2x - 24 = 0

factoring ___ (x - 6)(x + 4) = 0


HINT : logarithms are not defined for negative arguments , so only one of the factored solutions is correct